# A Novel Framework Based on Variational Quantum Algorithms: Revolutionizing Image Classification

### 13 Dec 2023 v1

@@ Abstract
Image classification is a crucial task in machine learning. In recent years, this field has witnessed rapid development, with a series of image classification models being proposed and achieving state-of-the-art (SOTA) results. Parallelly, with the advancement of quantum technologies, quantum machine learning has attracted a lot of interest. In particular, a class of algorithms known as variational quantum algorithms (VQAs) has been extensively studied to improve the performance of classical machine learning. In this paper, we propose a novel image classification framework using VQAs. The major advantage of our framework is the elimination of the need for the global pooling operation typically performed at the end of classical image classification models. While global pooling can help to reduce computational complexity, it often results in a significant loss of information. By removing the global pooling module before the output layer, our approach allows for effectively capturing more discriminative features and fine-grained details in images, leading to improved classification performance. Moreover, employing VQAs enables our framework to have fewer parameters compared to the classical framework, even in the absence of global pooling, which makes it more advantageous in preventing overfitting. We apply our method to different SOTA image classification models and demonstrate the superiority of the proposed quantum architecture over its classical counterpart through a series of experiments on public datasets.
@@ 1 Introduction
In contrast to the above methods, our work employs quantum machine learning algorithms to address the issue of global pooling for the task of image classification. Quantum machine learning has the potential to revolutionize the area of image classification as quantum algorithms can process large datasets of images more efficiently than classical algorithms, leading to faster and more accurate classification [dang2018image, kerenidis2019quantum, mari2020transfer, henderson2020quanvolutional, li2020quantum, liu2021hybrid, henderson2021methods, houssein2021hybrid, sagingalieva2022hyperparameter, hur2022quantum, senokosov2023quantum, melnikov2023quantum]. In this paper, we propose a novel image classification framework based on variational quantum algorithms (VQAs) [mcclean2016theory, bharti2022noisy, cerezo2021variational] in the context of quantum machine learning. VQAs are a class of quantum algorithms that utilize classical computers to optimize a parameterized quantum circuit. Due to their hybrid quantum-classical approach that has potential noise resilience [mcclean2016theory], VQAs have been considered as one of the leading candidates to achieve application-oriented quantum computational advantage on NIhyperrefSQ devices.

### 06 Mar 2024 v2

# A Novel Image Classification Framework Based on Variational Quantum Algorithms

@@ Abstract
Image classification is a crucial task in machine learning with widespread practical applications. The existing classical framework for image classification typically utilizes a global pooling operation at the end of the network to reduce computational complexity and mitigate overfitting. However, this operation often results in a significant loss of information, which can affect the performance of classification models. To overcome this limitation, we introduce a novel image classification framework that leverages variational quantum algorithms (VQAs)—hybrid approaches combining quantum and classical computing paradigms within quantum machine learning. The major advantage of our framework is the elimination of the need for the global pooling operation at the end of the network. In this way, our approach preserves more discriminative features and fine-grained details in the images, which enhances classification performance. Additionally, employing VQAs enables our framework to have fewer parameters than the classical framework, even in the absence of global pooling, which makes it more advantageous in preventing overfitting. We apply our method to different state-of-the-art image classification models and demonstrate the superiority of the proposed quantum architecture over its classical counterpart through a series of experiments on public datasets. Our experiments show that the proposed quantum framework achieves up to a 9.21% increase in accuracy and up to a 15.79% improvement in F1 score, compared to the classical framework.
@@ 1 Introduction
In contrast to the above methods, our work employs quantum machine learning algorithms to address the issue of global pooling for the task of image classification. Quantum machine learning has the potential to revolutionize the area of image classification as quantum algorithms can process large datasets of images more efficiently than classical algorithms, leading to faster and more accurate classification [dang2018image, kerenidis2019quantum, mari2020transfer, henderson2020quanvolutional, li2020quantum, liu2021hybrid, henderson2021methods, houssein2021hybrid, sagingalieva2022hyperparameter, hur2022quantum, senokosov2023quantum, melnikov2023quantum]. In this paper, we propose a novel image classification framework based on variational quantum algorithms (VQAs) [mcclean2016theory, bharti2022noisy, cerezo2021variational] in the context of quantum machine learning. VQAs are a class of quantum algorithms that utilize classical computers to optimize a parameterized quantum circuit. Due to their hybrid quantum-classical approach that has potential noise resilience [mcclean2016theory], VQAs have been considered as one of the leading candidates to achieve application-oriented quantum computational advantage on NISQ devices.